# Explainable Metrics for the Assessment of Neurodegenerative Diseases through Handwriting Analysis

### v3

@@ IV-A3 PD vs PDM
Among the 54 task-agnostic metrics, only 6 are significant ($p\leq 0.05$) in comparing PD and PDM participants for at least one task. We only found significant metrics for the Spiral tasks here.
To distinguish PD from other parkinsonism, significant variations can only be found in the Spiral tasks, motor exercises designed for PD assessment. We observe few significant metrics for the Spiral tasks, mostly amplitudes of speed and acceleration, the best being for the vertical velocity. The Spiral tasks reveal a difference in speed between PD and PDM participants, the latest being slower, as shown in Figure 11.
Figure 11: Boxplots of some the best metrics extracted for PD vs PDM for all the spiral tasks combined.
Figure 12: Boxplots of the best point-specific metrics for all groups of participants.
@@ IV-B Task-Specific Results
After presenting the results of the task-agnostic metrics, we focus on the task-specific ones, which are crafted especially for certain tasks or groups of tasks.
@@ IV-B1 Point Tasks Results
Figure 12 shows all metrics with their significance. None of the metrics are consistently significant across all tasks, but the radius average and radius standard deviation metrics exhibit the same behavior across all tasks: metrics from AD participants are higher than in the other groups. We observe a slightly higher radius average for PD participants compared to CTL, only significant when considering all point tasks, showing that the Point tasks capture only little of the expected tremors.
@@ IV-B2 Spiral Tasks Results
Figure 13 shows the boxplots of the spiral-specific metrics that were significant for at least one of the tasks. All the metrics exhibit a similar behavior across all tasks: PDM and AD participants take more time to draw their spirals than PD and CTL, AD and CTL participants draw more loops, which gives a lower time per loop (Fundamental period of the spiral) for PDM and AD participants. Subjects were not instructed to provide a minimum number of loops, therefore interpretation of task duration is difficult. The reduced number of loops in the PD cohort may have been an energetically rationale choice as the task is particularly difficult with persons with impaired dexterity. As PDM represents a mixed disease cohort, few conclusions relative to PD can be drawn. It is doubtful differences are due to cognitive impairment as overall MOCA performance was similar between these groups.
TABLE V: $p$-values and AUC of the specific metrics for all spiral tasks significant for at least one comparison. Significant p-values and AUC over 0.75 are bolded.
Spirals Metric AD vs CTL CTL vs PD PD vs PDM
@@
Figure 13: Boxplots of the best spiral-specific metrics, for all groups of participants.
@@ IV-B3 Numbers Task Results
TABLE VI: $p$-values and AUC of the significant metrics specific to the numbers task. Significant p-values and AUC over 0.75 are bolded.
Metric AD vs CTL CTL vs PD PD vs PDM
@@
Table VI shows the Numbers-specific metrics' results, studying mostly the variations in time and dimensions between successive answers in the Numbers Task. Of the few metrics that show a significant difference between groups of participants, the time taken to answer is the first metric differentiating between AD and CTL participants, AD participants being slower. In a similar fashion, AD participants are significantly slower over time as compared to CTL participants, an effect of the increased fatigue over the various answers. Finally, we observe that both AD and PD participants tend to input more and more pressure over time compared to CTL participants, once again a potential fatigue indicator.
@@ IV-B4 Writing Tasks Results
Figure 14: Boxplots of the best writing specific metrics for all groups of participants.
In this section, we present the results for the writing tasks's specific metrics. As shown in Figure 14, AD participants take a much longer Average Time Between Writing than all other participants across all tasks, revealing a longer hesitation and intermitent pausing times, between words. CTL participants generally have lower variations in the pressure they use than ND participants, across all tasks, a potential indicator of more stability in steadiness in the writing. PD participants use less and less pressure as they write, in contrast to the AD and CTL participants, and they also use a larger area to write, significant only in the Copy Text Task, which correlates with previous findings about PD participants writing larger than others [1].
TABLE VII: $p$-values and AUC of the specific metrics for all significant writing tasks. Significant p-values and AUC over 0.75 are bolded.
Writing Metric AD vs CTL CTL vs PD PD vs PDM
@@ IV-C Classification Results
TABLE VIII: Classification results for pairwise group comparisons, using the Bagging model. The table reports experimental results in terms of accuracy (ACC), F1-score (F1), and AUC for each comparison. The best results for each metric for each group comparison are indicated in bold, while the second-best are italicized.
Task AD vs PD PD vs PDM AD vs CTL AD vs PDM PD vs CTL PDM vs CTL
@@
When comparing AD against PD, the Copy Image and the Point Dominant tasks exhibit the best performance in terms of accuracy (91) and F1-Score (91). The Copy Image task involves visual perception and fine motor skills, the latter of which is particularly impaired in PD. Moreover, the Point Non-Dominant task reports the highest AUC score (94), indicating its efficacy in discriminating between the two conditions [53]. This is intriguing as mirror movement or mirroring dystonia have been described in Task Specific Dystonia and other movement disorders as a way to unmask and characterize latent dystonia metrics [54]; in our PD cohort, this task may be abnormal in PD subjects who are prone to the development of dystonic metrics - a finding generally thought to be absent in AD.
@@
For the comparison between AD and CTL, the Draw Clock task achieves the highest accuracy (87) and F1 score (87). This task assesses the ability to draw a clock with two hands correctly placed. This complex task includes a wide set of skills, including adequate function of spatial awareness, executive planning, error detection, motor coordination and execution as well as language processing and numerical processing. This task is often used to screen for intra-cortical and subcortical brain network deficiencies, observed notably in individuals with early-stage dementia. The high discriminability of this task could be attributed to cognitive impairment present in the AD group, as indicated by their lower MoCA score compared to other groups. Additionally, the Copy Text and Point Sustained tasks yield the highest AUC, indicating their utility in distinguishing between AD and CTL groups.
When contrasting AD with PDM, the highest accuracy (83) and the F1 score (83) are achieved in the Point Sustained task. Given the attentional deficits commonly observed in AD, this task's effectiveness in discriminating between AD and CTL and PDM groups is plausible.
@@
IV-D Correlation of the metrics with UPDRS-III and MoCa scores
To further explore the clinical relevance of the extracted handwriting metrics, we computed the Pearson correlation coefficient between each of the 54 task-agnostic metrics and the MoCA score [34], which evaluates global cognitive functioning, and the motor section of the Unified Parkinson's Disease Rating Scale (UPDRS-III) [35]. Only 13 out of 54 were significantly correlated with UPDRS-III scores ($p<0.05$), but none of them had $p$-values below 0.001, suggesting limited direct association between our handwriting metrics and motor symptom severity as assessed by UPDRS-III. However, a categorical examination of each subdomain of the UPDRS-III reveal a great variation of correlations depending on the domains, as show in Table IX. Asexpected, the tremor components are much more mesurable through handwriting than the postural and gait ones. The complete list of metrics with their individual significance is available in the appendix.
TABLE IX: Number of metrics over the 54 task-agnostic metrics produced being correlated with MoCa score, UPDRS-III score and its sub-components.
Score $p<0.05$ $p<0.01$ $p<0.001$ $p<0.0001$
MoCa [34] 33 22 15 9
UPDRS-III [35] 13 10 0 0
Rigidity 4 0 0 0
Upper bradykinesia 0 0 0 0
Lower bradykinesia 22 0 0 0
Arising from chair 2 0 0 0
Gait 20 4 0 0
Freeze of gait 8 2 2 2
Postural stability 0 0 0 0
Posture 1 1 0 0
@@ IV-D Correlation of the metrics with UPDRS-III and MoCa scores
Kinetic tremor 19 5 0 0
Postural Tremor 15 6 0 0
Resting tremor 27 23 13 2
Global tremor 30 23 12 4
Non-tremor components 0 0 0 0
In contrast, a substantial number of metrics (33 out of 54) were significantly correlated with MoCA scores ($p<0.05$), indicating a strong relationship between handwriting behavior and cognitive performance. Among these, 9 metrics exhibited highly significant correlations with $p$-values below 0.0001, underscoring the potential of handwriting-based metrics as sensitive indicators of cognitive impairment.
@@ V Discussion and Conclusion
We proposed and evaluated a large array of interpretable handwriting metrics designed to capture motor, cognitive, and behavioral markers from handwriting signals. These metrics were categorized as either task-agnostic (general signal descriptors applicable across tasks) or task-specific (metrics tailored to the expected behavior in particular task types). Our aim was to assess the ability of these metrics to distinguish between ND subtypes and healthy controls using statistical analysis and binary classification experiments, while ensuring clinical interpretability and relevance.
V-A Task-Agnostic and Task-Specific Metrics: Findings and Limitations
Our analysis of task-agnostic metrics, those computed independently of the specific task performed, demonstrated that a substantial number (53 out of 54) significantly differentiated at least one group comparison (e.g., PD vs. control, AD vs. PDM). These metrics included speed, acceleration, angular velocity, entropy, and timing-related metrics. Notably, these general metrics were consistent across task categories and aligned with known ND symptomatology. For example, reduced speed and increased entropy were particularly salient in AD participants, aligning with executive dysfunction and impaired motor planning.
However, the task-agnostic approach also revealed limitations. While general metrics captured high-level differences, they occasionally failed to detect hallmark disease-specific signs. Most notably, micrographia, a classic symptom of PD, was not captured in a statistically significant way. This suggests that some ND-specific metrics may only manifest under conditions that tightly control for task type, hand dominance, or cognitive load.
To address this limitation, we introduced task-specific metrics, tailored to the intent and structure of each task. These revealed more nuanced and disease-relevant patterns. AD participants showed longer in-air times, slower execution, and greater pressure variability, particularly during cognitively demanding tasks such as paragraph writing and arithmetic. PD participants drew fewer spiral loops, moved more slowly during motor tasks, and exhibited greater pressure variability, overlapping with patterns observed in the PDM group. While the PDM group was younger, its behavior often fell between the PD and control groups, suggesting that our metrics are relatively robust to age differences.
@@ V-B Classification Performance and Observed Limitations
In this study, we prioritized classification using interpretable metrics, incorporating task-specific and general metrics across handwriting tasks. Our primary goal was to develop a model that provides meaningful insights into the motor and cognitive impairments associated with neurodegenerative diseases, ensuring that the extracted metrics remain clinically interpretable. This approach allows for a better understanding of the underlying mechanisms driving classification decisions, which is crucial for real-world applications in clinical settings. However, deep learning-based approaches, such as those leveraging raw handwriting images or spectrogram representations of handwriting signals, offer a promising alternative. In a recent study [55], we explored non-interpretable methods by applying CNN and CNN-BLSTM models to spectrogram representations of handwriting signals for neurodegenerative disease classification. Our results demonstrated that classification performance varied across handwriting tasks and spectrogram channel combinations, with CNN consistently outperforming CNN-BLSTM. While the current study focused on interpretable metrics, future work could investigate hybrid approaches integrating deep learning techniques with interpretable models to enhance classification performance while maintaining clinical transparency. This would allow for the benefits of deep learning’s feature extraction capabilities while ensuring that the results remain interpretable and useful for clinical decision-making.
@@ V-C Clinical Applications and Implications
Handwriting analysis has strong potential as a non-invasive, low-cost, and accessible digital biomarker. It captures both motor and cognitive processes and can be easily deployed using commercial tablets or styluses. Clinically, such tools could support early screening, aid differential diagnosis, and monitor disease progression or response to treatment. The metrics developed in this study are interpretable and aligned with known clinical metrics, making them suitable for real-world use.
@@ V-D Discussion
While the proposed metrics and classification results offer compelling evidence for the utility of handwriting in distinguishing neurodegenerative conditions, there are several limitations that affect generalizability. First, the models developed in this study are trained on a dataset collected under tightly controlled conditions with a homogeneous set of equipment and task protocols. Their applicability to new populations, clinical settings, or hardware setups remains to be validated. This lack of generalization could be addressed through the use of pre-trained systems, such as neural networks which can learn more transferable representations, or through the use of more diverse handwriting datasets.
@@ V-E General Conclusion and Future Directions
This study introduced a diverse set of interpretable metrics for handwriting analysis and demonstrated their value in differentiating between neurodegenerative conditions using data from structured tablet-based tasks. The results validate handwriting as a meaningful digital biomarker and provide a strong foundation for future research and clinical translation.
@@ Appendix A Additional Tables
MoCa UPDRS Rigidity Upper Lower Arising Gait Freeze Postural Posture Gait and Kinetic Postural Resting Global Non

### v4

@@ IV-A3 PD vs PDM
Among the 54 task-agnostic features, only 6 are significant ($p\leq 0.05$) in comparing PD and PDM participants for at least one task. We only found significant features for the Spiral tasks here.
To distinguish PD from other parkinsonism, significant variations can only be found in the Spiral tasks, motor exercises designed for PD assessment. We observe few significant features for the Spiral tasks, mostly amplitudes of velocity and acceleration, the best being for the vertical velocity. The Spiral tasks reveal a difference in velocity between PD and PDM participants, the latest being slower, as shown in Figure 11.
Figure 11: Boxplots of some the best features extracted for PD vs PDM for all the spiral tasks combined.
Figure 12: Boxplots of the best point-specific features for all groups of participants.
@@ IV-B Task-Specific Results
After presenting the results of the task-agnostic features, we focus on the task-specific ones, which are crafted especially for certain tasks or groups of tasks.
@@ IV-B1 Point Tasks Results
Figure 12 shows all features with their significance. None of the features are consistently significant across all tasks, but the radius average and radius standard deviation features exhibit the same behavior across all tasks: features from AD* participants are higher than in the other groups. We observe a slightly higher radius average for PD participants compared to CTL, only significant when considering all point tasks, showing that the Point tasks capture only little of the expected tremors.
@@ IV-B2 Spiral Tasks Results
Figure 13 shows the boxplots of the spiral-specific features that were significant for at least one of the tasks. All the features exhibit a similar behavior across all tasks: PDM and AD* participants take more time to draw their spirals than PD and CTL, AD* and CTL participants draw more loops, which gives a lower time per loop (Fundamental period of the spiral) for PDM and AD* participants. Subjects were not instructed to provide a minimum number of loops, therefore interpretation of task duration is difficult. The reduced number of loops in the PD cohort may have been an energetically rationale choice as the task is particularly difficult with persons with impaired dexterity. As PDM represents a mixed disease cohort, few conclusions relative to PD can be drawn. It is doubtful differences are due to cognitive impairment as overall MOCA performance was similar between these groups.
TABLE V: $p$-values and AUC of the specific features for all spiral tasks significant for at least one comparison. Significant p-values and AUC over 0.75 are bolded.
Spirals Feature AD* vs CTL CTL vs PD PD vs PDM
@@
Figure 13: Boxplots of the best spiral-specific features, for all groups of participants.
@@ IV-B3 Numbers Task Results
TABLE VI: $p$-values and AUC of the significant features specific to the numbers task. Significant p-values and AUC over 0.75 are bolded.
Feature AD* vs CTL CTL vs PD PD vs PDM
@@
Table VI shows the Numbers-specific features' results, studying mostly the variations in time and dimensions between successive answers in the Numbers Task. Of the few features that show a significant difference between groups of participants, the time taken to answer is the first feature differentiating between AD* and CTL participants, AD* participants being slower. In a similar fashion, AD* participants are significantly slower over time as compared to CTL participants, an effect of the increased fatigue over the various answers. Finally, we observe that both AD* and PD participants tend to input more and more pressure over time compared to CTL participants, once again a potential fatigue indicator.
@@ IV-B4 Writing Tasks Results
Figure 14: Boxplots of the best writing specific features for all groups of participants.
In this section, we present the results for the writing tasks's specific features. As shown in Figure 14, AD* participants take a much longer Average Time Between Writing than all other participants across all tasks, revealing a longer hesitation and intermitent pausing times, between words. CTL participants generally have lower variations in the pressure they use than ND participants, across all tasks, a potential indicator of more stability in steadiness in the writing. PD participants use less and less pressure as they write, in contrast to the AD* and CTL participants, and they also use a larger area to write, significant only in the Copy Text Task, which correlates with previous findings about PD participants writing larger than others [1].
TABLE VII: $p$-values and AUC of the specific features for all significant writing tasks. Significant p-values and AUC over 0.75 are bolded.
Writing Feature AD* vs CTL CTL vs PD PD vs PDM
@@ IV-C Classification Results
TABLE VIII: Classification results for pairwise group comparisons, using the Bagging model. The table reports experimental results in terms of accuracy (ACC), F1-score (F1), and AUC for each comparison. The best results for each feature for each group comparison are indicated in bold, while the second-best are italicized.
Task AD* vs PD PD vs PDM AD* vs CTL AD* vs PDM PD vs CTL PDM vs CTL
@@
When comparing AD* against PD, the Copy Image and the Point Dominant tasks exhibit the best performance in terms of accuracy (91) and F1-Score (91). The Copy Image task involves visual perception and fine motor skills, the latter of which is particularly impaired in PD. Moreover, the Point Non-Dominant task reports the highest AUC score (94), indicating its efficacy in discriminating between the two conditions [54]. This is intriguing as mirror movement or mirroring dystonia have been described in Task Specific Dystonia and other movement disorders as a way to unmask and characterize latent dystonia features [55]; in our PD cohort, this task may be abnormal in PD subjects who are prone to the development of dystonic features - a finding generally thought to be absent in AD.
@@
For the comparison between AD* and CTL, the Draw Clock task achieves the highest accuracy (87) and F1 score (87). This task assesses the ability to draw a clock with two hands correctly placed. This complex task includes a wide set of skills, including adequate function of spatial awareness, executive planning, error detection, motor coordination and execution as well as language processing and numerical processing. This task is often used to screen for intra-cortical and subcortical brain network deficiencies, observed notably in individuals with early-stage dementia. The high discriminability of this task could be attributed to cognitive impairment present in the AD* group, as indicated by their lower MoCA score compared to other groups. Additionally, the Copy Text and Point Sustained tasks yield the highest AUC, indicating their utility in distinguishing between AD* and CTL groups.
When contrasting AD* with PDM, the highest accuracy (83) and the F1 score (83) are achieved in the Point Sustained task. Given the attentional deficits commonly observed in AD, this task's effectiveness in discriminating between AD* and CTL and PDM groups is plausible.
@@
IV-D Correlation of the features with UPDRS-III and MoCa scores
To further explore the clinical relevance of the extracted handwriting features, we measure the correlation between each of the 54 task-agnostic features and the MoCA score [35], which evaluates global cognitive functioning, and the motor section of the Unified Parkinson's Disease Rating Scale (UPDRS-III) [36], as well as each of the 14 subdomains of the UPDRS-III.
[…]
Then we verify the linearity condition between our features and the scores previously mentioned by fitting a linear and a quadratic function to each pair of features/scores, and perform a Nested model F-test over the residual sums of squares of both functions, to find out if adding the quadratic component makes a significant difference. Most features did not pass the linearity test with the various scores (3.4% to 16.3% had a p-value under 0.05, depending on the score at hand). As most features do not pass the linearity test, we use the Spearman correlation to measure the correlation between each score and each feature.
None of the features were significantly correlated with UPDRS-III scores ($p<0.05$), suggesting limited association between our handwriting features and motor symptom severity as assessed by UPDRS-III. However, UPDRS-III is a conglomerate of many sub-domains, for which some are not related to handwriting at all. A categorical examination of each subdomain of the UPDRS-III revealed a great variation of correlations depending on the domains, as show in Table IX. The complete list of features with their individual significance is available in the appendix.
TABLE IX: Number of features over the 54 task-agnostic features produced being correlated with MoCa score, UPDRS-III score and its sub-components.
Score p_value<p\_value<
$0.05$ $0.01$ $10^{-3}$ $10^{-4}$
MoCa [35] 30 18 11 9
UPDRS-III [36] 0 0 0 0
Rigidity 0 0 0 0
Upper bradykinesia 2 2 0 0
Lower bradykinesia 28 17 2 0
Arising from chair 15 4 1 0
Gait 19 0 0 0
Freeze of gait 16 6 0 0
Postural stability 7 3 0 0
Posture 0 0 0 0
@@ IV-D Correlation of the features with UPDRS-III and MoCa scores
Kinetic tremor 22 0 0 0
Postural Tremor 6 2 0 0
Resting tremor 6 0 0 0
Global tremor 33 3 0 0
Non-tremor components 1 0 0 0
In contrast, a substantial number of features (30 out of 54) were significantly correlated with MoCA scores ($p<0.05$), indicating a strong relationship between handwriting behavior and cognitive performance. Among these, 11 features exhibited highly significant correlations with $p$-values below 0.0001, underscoring the potential of handwriting-based features as sensitive indicators of cognitive impairment.
@@ V Discussion and Conclusion
We proposed and evaluated a large array of interpretable handwriting features designed to capture motor, cognitive, and behavioral markers from handwriting signals. These features were categorized as either task-agnostic (general signal descriptors applicable across tasks) or task-specific (features tailored to the expected behavior in particular task types). Our aim was to assess the ability of these features to distinguish between ND subtypes and healthy controls using statistical analysis and binary classification experiments, while ensuring clinical interpretability and relevance.
V-A Task-Agnostic and Task-Specific Features: Findings and Limitations
Our analysis of task-agnostic features, those computed independently of the specific task performed, demonstrated that a substantial number (53 out of 54) significantly differentiated at least one group comparison (e.g., PD vs. control, AD* vs. PDM). These features included velocity, acceleration, angular velocity, entropy, and timing-related features. Notably, these general features were consistent across task categories and aligned with known ND symptomatology. For example, reduced velocity and increased entropy were particularly salient in AD* participants, aligning with executive dysfunction and impaired motor planning.
However, the task-agnostic approach also revealed limitations. While general features captured high-level differences, they occasionally failed to detect hallmark disease-specific signs. Most notably, micrographia, a classic symptom of PD, was not captured in a statistically significant way. This suggests that some ND-specific features may only manifest under conditions that tightly control for task type, hand dominance, or cognitive load.
To address this limitation, we introduced task-specific features, tailored to the intent and structure of each task. These revealed more nuanced and disease-relevant patterns. AD* participants showed longer in-air times, slower execution, and greater pressure variability, particularly during cognitively demanding tasks such as paragraph writing and arithmetic. PD participants drew fewer spiral loops, moved more slowly during motor tasks, and exhibited greater pressure variability, overlapping with patterns observed in the PDM group. While the PDM group was younger, its behavior often fell between the PD and control groups, suggesting that our features are relatively robust to age differences.
@@ V-B Classification Performance and Observed Limitations
In this study, we prioritized classification using interpretable features, incorporating task-specific and general features across handwriting tasks. Our primary goal was to develop a model that provides meaningful insights into the motor and cognitive impairments associated with neurodegenerative diseases, ensuring that the extracted features remain clinically interpretable. This approach allows for a better understanding of the underlying mechanisms driving classification decisions, which is crucial for real-world applications in clinical settings. However, deep learning-based approaches, such as those leveraging raw handwriting images or spectrogram representations of handwriting signals, offer a promising alternative. In a recent study [56], we explored non-interpretable methods by applying CNN and CNN-BLSTM models to spectrogram representations of handwriting signals for neurodegenerative disease classification. Our results demonstrated that classification performance varied across handwriting tasks and spectrogram channel combinations, with CNN consistently outperforming CNN-BLSTM. While the current study focused on interpretable features, future work could investigate hybrid approaches integrating deep learning techniques with interpretable models to enhance classification performance while maintaining clinical transparency. This would allow for the benefits of deep learning’s feature extraction capabilities while ensuring that the results remain interpretable and useful for clinical decision-making.
@@ V-C Clinical Applications and Implications
Handwriting analysis has strong potential as a non-invasive, low-cost, and accessible digital biomarker. It captures both motor and cognitive processes and can be easily deployed using commercial tablets or styluses. Clinically, such tools could support early screening, aid differential diagnosis, and monitor disease progression or response to treatment. The features developed in this study are interpretable and aligned with known clinical features, making them suitable for real-world use.
@@ V-D Discussion
While the proposed features and classification results offer compelling evidence for the utility of handwriting in distinguishing neurodegenerative conditions, there are several limitations that affect generalizability. First, the models developed in this study are trained on a dataset collected under tightly controlled conditions with a homogeneous set of equipment and task protocols. Their applicability to new populations, clinical settings, or hardware setups remains to be validated. This lack of generalization could be addressed through the use of pre-trained systems, such as neural networks which can learn more transferable representations, or through the use of more diverse handwriting datasets.
@@ V-E General Conclusion and Future Directions
This study introduced a diverse set of interpretable features for handwriting analysis and demonstrated their value in differentiating between neurodegenerative conditions using data from structured tablet-based tasks. The results validate handwriting as a meaningful digital biomarker and provide a strong foundation for future research and clinical translation.
@@ Appendix A Additional Tables
Feature MoCa [35] UPDRS-III [36] Rigidity Upper Lower Arising Gait Freeze Postural Posture Gait and Kinetic Postural Resting Global Non-tremor